Vic is now 24 years older than his brother. in six years, he will be twice as older as his brother's age. What are their present ages
Answer by macston(5194)

V=Vic's age; B= brother's age
V=B+24 years; V+6years=2(B+6years)
(B+24 years)+6 years=2(B+6 years)
B+30 years=2B+12 years Subtract 12 from both sides
B+30 years-12 years=2B+12 years-12 years
B+18 years=2B Subtract B from both sides
B+18 years-B=2B-B
18=B ANSWER 1 Brother's age=18 years
V=B+24
V=18+24=42 ANSWER 2 Vic's age=42 years
CHECK:
In six years he will be twice as old as his brother.
V+6= 2(B+6)
42+6=2(28+6)
48=2(24)
48=48